# Exact Quantum Capacity of Decohering Channels in Arbitrary Dimensions

### v1

@@ Abstract
We derive an exact analytical expression for the quantum capacity of a broad class of decohering channels of the form $\Lambda(\rho)=(1-x)\rho+xD(\rho),$ where $D(\rho)$ denotes different forms of decoherence of the density matrix $\rho$. Unlike previous studies restricted to qubit systems or asymptotic bounds, our result holds for arbitrary finite-dimensional Hilbert spaces. We show that these channels are degradable for all $x$, leading to a closed-form, single-letter capacity formula that does not require regularization. Besides full decoherence, we extend our analysis to block-decohering channels, where decoherence occurs within orthogonal or overlapping subspaces. This family interpolates between identity channels and classical-quantum channels. Our findings clarify the role of coherence in quantum communication and provide rare exact benchmarks in the study of quantum channel capacities.
@@ 1 Introduction
Decoherence, the loss of quantum coherence due to environmental interactions, lies at the heart of quantum physics, governing the transition from quantum to classical behavior and imposing fundamental limits on quantum technologies [1, 2, 3]. It manifests as dephasing (phase randomization), amplitude damping (energy loss), or generalized noise, destroying superposition and entanglement both of which are essential for quantum computation , communication, and sensing[4, 5, 6]. While decoherence explains emergent classicality in macroscopic systems [7, 8, 2], its mitigation is critical for scalable quantum devices, where noise can destroy fragile quantum states. Understanding its precise impact on quantum information processing, particularly through the lens of channel capacities, remains a central challenge due to the superadditivity of coherent information [9, 10, 11, 12] and the need for regularization [13, 14, 15] in capacity calculations.
In this work, we address this challenge by deriving exact quantum capacities for a broad class of decohering channels, where noise is dominated by dephasing. While there are few channels for which quantum capacity is exactly known, like the bit flip channel [15] or the erasure channel [14], we identify a class of operationally important quantum channels, of the form
@@
where $D(\rho)$ induces complete dephasing. Understandably, the capacity of these channels reduce to zero, when decoherence is complete at $x=1$. We then obtain the exact quantum capacity for family of channels of the form
@@
where now block-decohrence [16, 17, 18, 19, 20, 21, 22, 23] occurs between orthogonal subspaces of dimension $k$ and is preserved in each subspace.
The surprize is when decoherence acts between subspaces rather than individual states: even under maximal noise ($x=1$), the channel retains non-zero capacity, if coherence persists within subspaces. This demonstrates that subspace-localized coherence, a form of "protected" quantum information, can defy complete erasure. Our explicit capacity formula quantifies this effect, and shows how decoherence-free subspaces [24] contribute in the capacity of this channel. These results establish rigorous benchmarks for fault-tolerant protocols and reveal hidden resources in partially coherent systems. Our results bridge foundational questions about the quantum-classical divide with practical constraints in quantum engineering, and offers tools to optimize information transmission in quantum networks and quantum memories.
To go beyond this decoherence-free regime, we derive the exact capacity for a third class of quantum channels of the form
@@
where $W_{k}$ destroys coherence between overlapping subspaces and there is no decoherence-free subspace. Even in this case, we see that capacity remains non-zero in the limit of strong decoherence.
Our method relies on an explict proof of degradability [25, 26] of all these channels, through a method which can set a method for constructing large families of degradable channels.
The structure of this paper is as follows: In section 2, we collect some known facts about the quantum capacity of quantum information and describe why it is do difficult to obtain exact results for capacities. In 3, we first study the case of fully decohering channel $\Lambda$, where the dimension of each subspace is $1$. After proving its degradability property we proceed to calculate its quantum capacity. We then go on in section 4 to consider the case where decoherence occurs between orthogonal subspaces of dimension $k$. The proofs and methods of calculation are similar to the case of $k=1$ with slight differences. Finally in section 5, we study the case of decoherence among overlapping subspaces. The basic resuts are shown in figures (1) and (2). The paper is concluded with a discussion.
@@ 2 Preliminaries on quantum capacities
An important development in this direction was made by introducing the concepts of degradablity of quantum channels [25, 26].
@@
[25, 26] Consider a quantum channel ${\cal E}:A\to B$ and its complementary channel ${\cal E}^{c}:A\to E$. The channel ${\cal E}$ is said to be degradable if there exists a quantum channel $\mathcal{M}:B\to E$ such that $\mathcal{M}\circ{\cal E}(\rho)={\cal E}^{c}(\rho)$. Conversely it is called anti-degradable if there is a quantum channel $\mathcal{N}:B\to E$ such that $\mathcal{N}\circ{\cal E}^{c}(\rho)={\cal E}(\rho)$.
@@
What we will show in this letter is to prove that the decohering channels $\Lambda,\Lambda_{k}$ and $\Phi_{k}$ are all degradable for all values of $x$ in every dimension and use this property to find their capacity in closed form. The results are shown in figures (1) and (2).
3 The decohering channel
@@
where $x\in[0,1]$ and $D(\rho)=\sum_{i=1}^{d}\rho_{ii}|i\rangle\langle i|$ shows the decohered state of $\rho$. This channel shows the effect of decohering noise which happens with probability $x$. The first basic property of this channel, which will play a central role in proving the degradability of the channel is that
$D(\Lambda(\rho))=D(\rho).$ (3)
@@
where $a=\sqrt{x(1-x)}$. The crucial property is that $\Lambda^{c}(Y)$ depends only on the diagonal entries of $Y$, that is
$\Lambda^{c}(Y)=\Lambda^{c}(D(Y)).$ (7)
This leads to the second basic property of the channel, namely
$\Lambda^{c}(\Lambda(\rho))=\Lambda^{c}\big{(}\ D(\Lambda(\rho))\ \big{)}.$ (8)
We now use combine the two basic properties of this channel (3) and (7) to prove that it is degradable. The reasoning is quite simple. We write
$\Lambda^{c}(\Lambda(\rho))=\Lambda^{c}(D(\Lambda(\rho)))=\Lambda^{c}(D(\rho)))=\Lambda^{c}(\rho),$ (9)
@@
Note that this is entirely different from self-complementary channels [39], i.e. those for which $\Lambda^{c}=\Lambda$. These channels are both degradable and anti-degradable and hence have zero capacity. The degradability of the decohering channel $\Lambda$, has now important consequences, namely that its full quantum capacity is the same as its one-shot capacity $Q^{(1)}(\Lambda).$ Moreover, degradability also implies that the coherent information $J(\Lambda,\rho)$ is a concave function of $\rho$. These two results greatly facilitates the exceedingly difficult task of optimization in eq. (1). First we can bypass the almost impossible task of regularization over infinitely many uses of the channel. Second we use concavity of the coherent information to do the optimization in a direct analytical way. To do this, we note that due to the Schur-Horn theorem in matrix theory, the diagonal entries of any Hermitian matrix are majorized by its eigenvalues. Thus we have
@@
which imediately leads to
@@
$J_{c}(\Lambda,\rho)\leq J_{c}(\Lambda,\rho_{d}).$ (13)
To find the state which maximizes the coherent information, we invoke the covariance property of the channel, namely
@@
where $U_{x}=|0\rangle\langle 0|\oplus X$. This means that $J_{c}(\rho)=J_{c}(X\rho X^{\dagger})$ for any density matrix. We now note that for diagonal density matrix $\rho_{d}$, it holds that
@@
$J_{c}(\frac{1}{d}I_{d})=J_{c}(\frac{1}{d}\sum_{i=1}^{d}X^{i}\rho_{d}{X^{\dagger}}^{i})\geq\frac{1}{d}\sum_{i=1}^{d}J_{c}(X^{i}\rho_{d}{X^{\dagger}}^{i})=\frac{1}{d}\sum_{i=1}^{d}J_{c}(\rho_{d})=J_{c}(\rho_{d})$ (18)
@@
$Q(\Lambda_{x})=\log d+\frac{d-1}{d}x\log\frac{x}{d}+(1-x+\frac{x}{d})\log(1-x+\frac{x}{d}).$ (22)
@@
4 Quantum Capacity of Block-decohering channels
@@
which affects shifts among subspaces. We now use this symmetry to show that for an arbitrary diagonal matrix,
$\frac{1}{d}\sum_{i=1}^{k}\sum_{j=1}^{r}U_{ij}\rho U_{ij}^{\dagger}=\frac{I}{d},$ (30)
@@
$Q(\Lambda_{k})=S(\Lambda_{k}(\frac{I}{d}))-S(\Lambda_{k}^{c}(\frac{I}{d}))$ (31)
@@
5 Quantum Capacity of Weakly-Decohering Channels
The non-zero capacity of block-decohering channels in the previous section may be ascribed to the presence of decoherence-free subspaces, namely the $k-$ dimensional decoherence-free subspaces which are not affected by the decohering noise. As we will see in this section, in the absence of these decoherence-free subspaces, and provided that some degree of coherence exists, the channel can still have non-zero quantum capacity. In this way, we are considering a noise which is not so strong to anihillate all the off-diagonal elements. Rather, we assume an equal number of minor diagonals above and below the main diagonal are also kept intact while all the other off-diagonal elements vanish. As an example, for $d=6$, an example of such weak decoherence is the following:
@@ 5 Quantum Capacity of Weakly-Decohering Channels
There is a related concept of coherence, called $k-$ coherence [22, 19], which is somewhat more general than this, but we will not consider it here. The original notion of incoherence [40], which defines incoherent states as diagonal density matrices in a specific basis, has been generalized to multi-level or k-coherence[19]. In this generalized setting, a state $|\psi\rangle=\sum_{i=1}^{d}\psi_{i}|i\rangle$ is said to have coherence at level k, if exactly $k$ of the coefficients $\psi_{i}$ are non-zero. Thus an incoherent state has coherence at level $1$, and a state like $|\psi\rangle=a|0\rangle+b|1\rangle$ in a $d-$ dimensional Hilbert space has coherence at level $2$ and so on. The generalization to mixed states is done by defining the states with coherence level $k$ to be the convex combination of all pure states whose coherence level is less than or equal to $k$. The notion of weak-coherence that we adopt here is different in that it assigns an order to the basis states. For example in the example (), If for some physical reason, say the photon number in each state, we order the states as $|0\rangle,|1\rangle,|2\rangle,\cdots$, two states of the form $|0\rangle+|1\rangle$ and $|1\rangle+|2\rangle$ are equally affected by decoherence, but a state like $|1\rangle+|3\rangle$ completely loses its coherence.
In this section, we consider decohering channels which with a certain probability decohere the input state to level-$k$ coherent states. Thats, we consider channesl of the form
@@
$D(\Phi_{k}(\rho))=\Phi_{k}(D(\rho))=D(\rho).$ (35)
For the second property (7), we show that for the channel $\Phi_{k}$, the complementary channel $\Phi_{k}^{c}$ still depends on the diagonal elements of $\rho$. To see this, it is sufficient to note that the Kraus operators $I$, $A_{i}$ and the products $A_{i}A_{j}$ are all projectors on subspaces and hence extarct the diagonal elements of $\rho$, when inserted in equation (5) for deriving the complementary channel. Therefore we have
$\Phi_{k}^{c}(\rho)=\Phi_{k}^{c}(D(\rho)).$ (36)
@@
$\Phi_{k}^{c}(\Phi_{k}(\rho))=\Phi_{k}^{c}(D(\Phi_{k}(\rho)))=\Phi_{k}^{c}(D(\rho))=\Phi_{k}^{c}(\rho),$ (37)
which shows that the channel $\Phi_{k}$ is also degradable. Invoking Schur-Horn theorem as before, we need only search among the diagonal density matrices and noting the symmetry of the channel
$\Phi_{k}(X\rho X^{\dagger})=X\Phi_{k}(\rho)X^{\dagger},$ (38)
where $X$ is the shift operator on the whole space, i.e. $X=\sum_{i=1}^{d}|i+1\rangle\langle i|$, we find from the concave property of the coherent information that the maximum value is obtained at $\rho_{d}=\frac{I}{d}$. Thus we have
$Q_{k}(\Phi)=I_{c}(\Phi_{k},\frac{I}{d}).$
We first calculate the following elements:
@@
$b(s)=\begin{cases}k-s&\text{if }\ \ 0\leq s\leq k-1,\\ 0&\text{if }\ \ k\leq s\leq d-k,\\ s+k-d&\text{if}\ \ d-k+1\leq s\leq d-1.\end{cases}$ (42)
@@
$\left[\begin{array}[]{c|c}1-x&\xi\sqrt{k}\bra{v_{0}}\\ \hline\cr\xi\sqrt{k}\ket{v_{0}}&xB\end{array}\right]$ (43)
where again $\xi=\sqrt{\frac{x(1-x)}{d}}$ and |$\ket{v_{0}}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\ket{i}$. and the matrix $B$ is a ciruclant matrix given by
@@
To find the quantum capacity of $\Phi_{k}$, we now need to diagonalize the matrix ${\cal M}:=\Phi_{k}^{c}(\frac{I}{d})$, as given in (43). We note that $|v_{0}\rangle$ is an eigenvector of $X$ and hence of $B$ with eigenvalue equal to $\lambda_{0}=\frac{k}{d}$. To find the spectrum of this matrix, we note that for any eigenvector $|v_{m\neq 0}\rangle$, the vector
@@
As a result, the set of eigenvalues of ${\cal M}=\Phi^{c}(\frac{I}{d})$ is:
@@
$Q(\Phi_{k})=I_{c}\left(\Phi,\frac{I}{d}\right)=\log{d}+\sum_{i=1}^{d-1}x\lambda_{i}\log(x\lambda_{i})+\left[\frac{kx}{d}+(1-x)\right]\log{\left[\frac{kx}{d}+(1-x)\right]}$ )
Figure 2: Coherent information $I_{c}\bigl{(}\Phi,\tfrac{I}{d}\bigr{)}$ versus $x$ for $d=12$ and $k\in\{1,2,3,4,6\}$.
As special cases we note that when $k=1$, this capacity reduces to that of (22), as expected. Also for $x=0$, or when $k=d$, for which the channel is the identity channel, the capacity is given by $Q(\Phi_{k})=\log_{2}d.$ Note that for the strongest level of noise $(x=1)$, the quantum capacity is no longer given by $\log_{2}k$. The capacities of these channels for $d=12$ and for various degrees of decoherence, determined by $k$ are ploted in figure (2). It is worth nothing that even at full decoherence, the channel has non-zero capacity, despite having no decoherence-free subspace.

### v2

@@ Abstract
We derive exact analytical expressions for the quantum capacity of a broad class of decohering quantum channels of the form $\Lambda(\rho)=(1-x)\rho+xD(\rho)$, where $D(\rho)$ represents a structured decoherence process. These channels are shown to be degradable for all noise parameters and in arbitrary dimensions, yielding closed-form, single-letter capacity formulas. Our analysis includes fully decohering, block-decohering, and weakly decohering channels, the latter involving coherence preservation within overlapping subspaces. Surprisingly, even under maximal decoherence, the channel may retain nonzero capacity due to residual coherence structure. These results provide quantitative role for decoherence-free and partially coherent subspaces in preserving quantum information, offering guidance for encoding strategies in quantum memories and fault-tolerant quantum communication systems.
@@ 1 Introduction
The capacity of a quantum channel to transmit information is a central figure of merit in quantum communication and quantum information processing [1, 2, 3]. Yet, despite its conceptual clarity, calculating the quantum capacity remains notoriously difficult, even for seemingly simple channels. This difficulty stems from the need to optimize the coherent information over multiple uses of the channel and the lack of additivity in general [4, 5, 6], rendering most capacities intractable. Exact capacity formulas are known only for a handful of special cases, such as the erasure channel [7], amplitude damping in limited regimes in low dimensions [8, 9, 10], and a few qubit Pauli channels [11]. In higher-dimensional systems, exact results are almost nonexistent [12, 13]. Even lower and upper bounds can be difficult [8, 14, 15, 16, 17, 18, 19, 20] to compute due to nonconvexities in the coherent information. Thus, any exact result in this domain not only deepens our theoretical understanding but also offers valuable benchmarks for real-world implementations. In this Letter, we present a broad family of quantum channels for which the quantum capacity can be computed exactly and in closed form. These channels take the form $\Lambda(\rho)=(1-x)\rho+xD(\rho)$, where $D(\rho)$ models structured decoherence [21, 22, 23, 24, 25]. Remarkably, we prove that these channels are degradable in all dimensions and for all values of the noise parameter x. This includes fully decohering channels, block-decohering models that preserve coherence within subspaces, and weakly decohering channels where coherence survives across overlapping modes. We show that even under strong noise $(x=1)$, these channels may retain nonzero quantum capacity—provided certain coherence structures are preserved [26, 27, 28, 29, 30, 25]. This finding challenges the common intuition that strong decoherence always implies complete quantum information loss.
We first consider the operationally important quantum channels, of the form
@@
where $D(\rho)$ induces complete dephasing, then we study channels of the form
@@
where now block-decoherence [31, 32, 33, 25, 34, 35, 36, 37] occurs between orthogonal subspaces of dimension $k$ and coherence is preserved in each subspace. The surprise is when decoherence acts between subspaces rather than individual states: even under maximal noise ($x=1$), the channel retains non-zero capacity, if coherence persists within subspaces. This demonstrates that subspace-localized coherence, a form of "protected" quantum information, can defy complete erasure. Our explicit capacity formula quantifies this effect, and shows how decoherence-free subspaces [38] contribute in the capacity of this channel. To go beyond these types of direct sum channels, we proceed to derive the exact capacity for a third class of quantum channels of the form
@@
where $W_{k}$ destroys coherence between overlapping subspaces in such a way that there is no decoherence-free subspace. Even in this case, we see that capacity remains non-zero in the limit of strong decoherence.
Our method relies on an explicit proof of degradability [8, 39] of all these channels, through a method which can set a new method for constructing large families of degradable channels.
The structure of this paper is as follows: In section 2, we collect some known facts about the quantum capacity of quantum information and describe why it is so difficult to obtain exact results for capacities. In 3, we first study the case of fully decohering channel $\Lambda$, where the dimension of each subspace is $1$. After proving its degradability property we proceed to calculate its quantum capacity. We then go on in section 4 to consider the case where decoherence occurs between orthogonal subspaces of dimension $k$. The proofs and methods of calculation are similar to the case of $k=1$ with slight differences. Finally in section 5, we study the case of decoherence among overlapping subspaces. The basic results are shown in figures (1) and (2). The paper is concluded with a discussion.
@@ 2 Preliminaries on quantum capacities
An important development in this direction was made by introducing the concepts of degradability of quantum channels [8, 39].
@@
[8, 39] Consider a quantum channel ${\cal E}:A\to B$ and its complementary channel ${\cal E}^{c}:A\to E$, where $A$ and $B$ denote the domain and target systems and $E$ the auxiliary system playing the role of environment. The channel ${\cal E}$ is said to be degradable if there exists a quantum channel $\mathcal{M}:B\to E$ such that $\mathcal{M}\circ{\cal E}={\cal E}^{c}$. Conversely it is called anti-degradable if there is a quantum channel $\mathcal{N}:E\to B$ such that $\mathcal{N}\circ{\cal E}^{c}={\cal E}$.
@@
What we will show in this paper is to prove that the decohering channels $\Lambda,\Lambda_{k}$ and $\Phi_{k}$ are all degradable for all values of $x$ in every dimension and use this property to find their capacities in closed form. The results are shown in figures (1) and (2).
3 The fully-decohering channel $\Lambda$
@@
where $x\in[0,1]$ and $D(\rho)=\sum_{i=1}^{d}\rho_{ii}|i\rangle\langle i|$ shows the fully decohered state of $\rho$. This channel shows the effect of decohering noise which happens with probability $x$. The first basic property of this channel, which will play a central role in proving the degradability of the channel is that for all input states $\rho$, $D(\Lambda(\rho))=D(\rho).$ or simply
$D\circ\Lambda=D.$ (3)
@@
where $a=\sqrt{x(1-x)}$. The second crucial property is that $\Lambda^{c}(Y)$ depends only on the diagonal entries of $Y$, that is, for any input state $\rho$
$\Lambda^{c}=\Lambda^{c}\circ D.$ (7)
The combination of (3) and (7)leads to the following chain of identities
$\Lambda^{c}\circ\Lambda=(\Lambda^{c}\circ D)\circ\Lambda=\Lambda^{c}\circ(D\circ\Lambda)=\Lambda^{c}\circ D=\Lambda^{c},$ (8)
@@
Note that this is entirely different from self-complementary channels [57], i.e. those for which $\Lambda^{c}=\Lambda$. These channels are both degradable and anti-degradable and hence have zero capacity. The degradability of the decohering channel $\Lambda$, has now important consequences, namely that its full quantum capacity is the same as its one-shot capacity $Q^{(1)}(\Lambda).$ Moreover, degradability also implies that the coherent information $J(\Lambda,\rho)$ is a concave function of $\rho$. These two results greatly facilitate the exceedingly difficult task of optimization in eq. (1). First we can bypass the almost impossible task of regularization over infinitely many uses of the channel. Second we use concavity of the coherent information to do the optimization in a direct analytical way. To do this, we note that due to the Schur-Horn theorem in matrix theory, the diagonal entries of any Hermitian matrix are majorized by its eigenvalues. Thus we have
@@
which immediately leads to
@@
$I_{c}(\Lambda,\rho)\leq I_{c}(\Lambda,\rho_{d}).$ (12)
[…]
@@
where $U_{x}=|0\rangle\langle 0|\oplus X$. This means that $I_{c}(\Lambda,\rho)=I_{c}(\Lambda,X\rho X^{\dagger})$ for any density matrix. We now note that for diagonal density matrix $\rho_{d}$, it holds that
@@
$I_{c}(\Lambda,\rho_{d})=\frac{1}{d}\sum_{i=1}^{d}I_{c}(\Lambda,\rho_{d})=\frac{1}{d}\sum_{i=1}^{d}I_{c}(\Lambda,X^{i}\rho_{d}{X^{\dagger}}^{i})\leq I_{c}(\Lambda,\frac{1}{d}\sum_{i=1}^{d}X^{i}\rho_{d}{X^{i}}^{\dagger})=I_{c}(\Lambda,\frac{1}{d}I_{d})$ (17)
@@
$Q(\Lambda)=\log d+\frac{d-1}{d}x\log\frac{x}{d}+(1-x+\frac{x}{d})\log(1-x+\frac{x}{d}).$ (21)
@@
4 The Block-decohering channel $\Lambda_{k}$
@@
which affects shifts among subspaces. We now use this symmetry to show that for an arbitrary diagonal matrix $\rho_{d}$,
$\frac{1}{d}\sum_{i=1}^{k}\sum_{j=1}^{r}U_{ij}\rho_{d}U_{ij}^{\dagger}=\frac{I}{d},$ (29)
@@
$\max_{\rho}I_{c}(\Lambda_{k},\rho)=S(\Lambda_{k}(\frac{I}{d}))-S(\Lambda_{k}^{c}(\frac{I}{d}))$ (30)
@@
5 The Weakly-Decohering Channel $\Phi_{k}$
The non-zero capacity of block-decohering channels in the previous section may be attributed to the presence of decoherence-free subspaces, namely the $k-$ dimensional decoherence-free subspaces which are not affected by the decohering noise. As we will see in this section, in the absence of these decoherence-free subspaces, and provided that some degree of coherence exists, the channel can still have non-zero quantum capacity. In this way, we are considering a noise which is not so strong to annihilate all the off-diagonal elements. Rather, we assume an equal number of minor diagonals above and below the main diagonal are also kept intact while all the other off-diagonal elements vanish. As an example, for $d=6$, an example of such weak decoherence is the following:
@@ 5 The Weakly-Decohering Channel $\Phi_{k}$
Note that for simplicity of calculation we assume a periodic boundary condition on the states, i.e. $|d+1\rangle\equiv|1\rangle$, that is why in this example the two elements in the corner are non-vanishing. There is a related concept of coherence, called $k-$ coherence [36, 25], which is somewhat more general than this, but we will not consider it here. The original notion of incoherence [58], which defines incoherent states as diagonal density matrices in a specific basis, has been generalized to multi-level or k-coherence [25]. In this generalized setting, a state $|\psi\rangle=\sum_{i=1}^{d}\psi_{i}|i\rangle$ is said to have coherence at level k, if exactly $k$ of the coefficients $\psi_{i}$ are non-zero. Thus an incoherent state has coherence at level $1$, and a state like $|\psi\rangle=a|0\rangle+b|1\rangle$ in a $d-$ dimensional Hilbert space has coherence at level $2$ and so on. The generalization to mixed states is done by defining the states with coherence level $k$ to be the convex combination of all pure states whose coherence level is less than or equal to $k$. The notion of weak-coherence that we adopt here is different in that it assigns an order to the basis states. For example in a $d-$ dimensional space, if for some physical reason, say the photon number in each state, we order the states as $|0\rangle,|1\rangle,|2\rangle,\cdots$, two states of the form $|0\rangle+|1\rangle$ and $|1\rangle+|2\rangle$ are equally affected by decoherence, but a state like $|1\rangle+|3\rangle$ completely loses its coherence.
In this section, we consider decohering channels which with a certain probability decohere the input state to level-$k$ coherent states of the type we defined above. That is, we consider channels of the form
@@
$D\circ\Phi_{k}=\Phi_{k}\circ D=D.$ (34)
For the second property (7), we show that for the channel $\Phi_{k}$, the complementary channel $\Phi_{k}^{c}$ still depends on the diagonal elements of $\rho$. To see this, it is sufficient to note that the Kraus operators $I$, $A_{i}$ and the products $A_{i}A_{j}$ are all projectors on subspaces and hence extract the diagonal elements of $\rho$, when inserted in equation (5) for deriving the complementary channel. Therefore we have
$\Phi_{k}^{c}=\Phi_{k}^{c}\circ D.$ (35)
@@
$\Phi_{k}^{c}\circ\Phi_{k}=(\Phi_{k}^{c}\circ D)\circ\Phi_{k}=\Phi_{k}^{c}\circ(D\circ\Phi_{k})=\Phi_{k}^{c}\circ D=\Phi_{k}$ (36)
which shows that the channel $\Phi_{k}$ is also degradable. Invoking Schur-Horn theorem as before, we need only search among the diagonal density matrices. We also note from (33) that $XA_{0}X^{\dagger}=A_{0},$ and $XA_{i}X^{\dagger}=A_{i+1}$, which show that the channel is covariant under shift operator $X$, that is
$\Phi_{k}(X\rho X^{\dagger})=X\Phi_{k}(\rho)X^{\dagger}.$ (37)
Combining these, we find from the concave property of the coherent information that the maximum value is obtained at $\rho_{d}=\frac{I}{d}$. Thus we have
$Q(\Phi_{k})=\max_{\rho_{d}}=I_{c}(\Phi_{k},\frac{I}{d}).$
To proceed we define
${\cal M}:=\Phi_{k}(\frac{I}{d})$
and calculate the following elements:
@@
$b(s):=\begin{cases}k-s&\text{if }\ \ 0\leq s\leq k-1,\\ 0&\text{if }\ \ k\leq s\leq d-k,\\ s+k-d&\text{if}\ \ d-k+1\leq s\leq d-1.\end{cases}$ (41)
@@
${\cal M}=\left[\begin{array}[]{c|c}1-x&\xi\sqrt{k}\bra{v_{0}}\\ \hline\cr\xi\sqrt{k}\ket{v_{0}}&xB\end{array}\right]$ (42)
where again $\xi=\sqrt{\frac{x(1-x)}{d}}$ and |$\ket{v_{0}}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\ket{i}$. and the matrix $B$ is a circulant matrix given by
@@
To find the quantum capacity of $\Phi_{k}$, we now need to diagonalize the matrix ${\cal M}\equiv\Phi_{k}^{c}(\frac{I}{d})$, as given in (42). We note that $|v_{0}\rangle$ is an eigenvector of $X$ and hence of $B$ with eigenvalue equal to $\lambda_{0}=\frac{k}{d}$. To find the spectrum of this matrix, we note that for any eigenvector $|v_{m\neq 0}\rangle$, the vector
@@
As a result, the set of eigenvalues of ${\cal M}=\Phi_{k}^{c}(\frac{I}{d})$ is:
@@
$Q(\Phi_{k})=I_{c}\left(\Phi_{k},\frac{I}{d}\right)=\log{d}+\sum_{i=1}^{d-1}x\lambda_{i}\log(x\lambda_{i})+\left[\frac{kx}{d}+(1-x)\right]\log{\left[\frac{kx}{d}+(1-x)\right]}$ )
Figure 2: Coherent information $\lambda\bigl{(}\Phi,\tfrac{I}{d}\bigr{)}$ versus $x$ for $d=12$ and $k\in\{1,2,3,4,6\}$.
As special cases we note that when $k=1$, this capacity reduces to that of (21), as expected. Also for $x=0$, or when $k=d$, for which the channel is the identity channel, the capacity is given by $Q(\Phi_{k})=\log_{2}d.$ Note that for the strongest level of noise $(x=1)$, the quantum capacity is no longer given by $\log_{2}k$. The capacities of these channels for $d=12$ and for various degrees of decoherence, determined by $k$ are plotted in figure (2). It is worth nothing that even at full decoherence, the channel has non-zero capacity, despite having no decoherence-free subspace.